# Hidden Markov modelling of spatio-temporal dynamics of measles in 1750–1850 Finland

### v1

@@ 1 Introduction
Describing, visualising, and modelling spatio-temporal panel data is often challenging, particularly in high-dimensional settings with numerous time points and spatial areas. Conditional autoregressive models (CAR) (Besag,, 1974) and their variants, such as intrinsic conditional autoregressive models (ICAR), are frequently employed to describe areal data. Although it is possible to extend CAR models to spatio-temporal settings (Clayton and Bernardinelli,, 1996), these interaction models can become computationally intensive with high-dimensional data where spatial and temporal effects are not separable (Knorr-Held,, 2000). Approximate methods like INLA (Rue et al.,, 2009) may be computationally efficient alternatives to asymptotically exact Bayesian estimation via Markov chain Monte Carlo (MCMC), but these methods typically impose some restrictions on the model structure and the number of hyperparameters (Bakka et al.,, 2018). The potential bias of these approximations is also challenging to quantify in specific cases, although in many typical settings the approximation bias has been found to be negligible (Ogden,, 2016; Rue et al.,, 2017). Furthermore, visualising and interpreting the results of these interaction models may be complex: with time-varying spatial effects and a large number of time points, a substantial number of static maps are required to convey information about the estimated (or observed) spatial dynamics.
@@
In life course research, hidden Markov models have been suggested as a probabilistic method to condense information of complex, multidomain life courses into a few interpretable latent states (Helske et al.,, 2018; Helske and Helske,, 2019). In these applications, given a common data generating process, individuals are assumed to be independent, having their own latent trajectories. In contrast, in spatio-temporal panel data, individuals share a spatial dependency structure which should be accounted for.
@@
In Amorós et al., (2020), hidden Markov models are utilised in a spatio-temporal setting to detect influenza outbreaks, and in Knorr-Held and Richardson, (2003), they are used to identify meningococcal disease incidences. Both studies use two states to identify the endemic and epidemic, or endemic and hyperendemic, phases of the infections but do not aim to deeply analyse the underlying dynamics of the diseases. Employing more states may aid in discovering latent phenomena related to the epidemics. Amorós et al., (2020) include the spatial dependence in the epidemic state as temporally independent ICAR components and omit it in the endemic state, whereas Knorr-Held and Richardson, (2003) account for the spatial dependency as ICAR in both states but consider it as time-invariant.
In this paper, we propose a Bayesian hidden Markov model with a state-dependent spatial correlation structure as an unsupervised learning method for decomposing the spatio-temporal variation in the data into relatively few interpretable latent components. We illustrate our approach by modelling the spatio-temporal occurrence of measles in 18th- and 19th-century Finland.
@@
Our model offers important insights into the historical measles epidemics in Finland. The extremely low population size coupled with hundreds of sparsely spaced small towns makes it challenging for epidemics to remain fully endemic (Keeling and Grenfell,, 1997). Thus, it is probable that fadeouts and reintroductions dominated the infection dynamics in Finland, implying that changes in the transmission sources could have strongly affected these dynamics. Finland and Sweden are separated by a large geographical barrier, the Gulf of Bothnia, and this, as well as the Russian border on the east, must have influenced the trade connections and the movement of people. The annexation of Finland to Russia, after a war in 1808–1809, resulted in the transfer of the capital from the western position of Turku to a more eastern location of Helsinki, increasing dramatically trade and social communication on the latter area, and drastically altering Finland’s connectivity to other countries (Ojala and Räihä,, 2017). We exploit our method to describe the reoccurring and presumably significantly foreign connection driven patterns of the epidemics in a compact yet informative way.
@@ 2.1 Data
Finland has a long history in collecting demographic and health-related data, with local parishes systematically maintaining records of baptisms, burials, and causes of death from 1749 onwards. Also earlier, albeit incomplete, records exist (Pitkänen,, 1977). Even though death diagnostics were symptom-based, the distinctive features of certain diseases allowed for relatively accurate diagnoses. One of those infections is measles, which we focus on. According to our data, measles accounted for approximately 2% of all deaths during our study period from 1750 to 1850. The population of Finland was roughly $500{,}000$ in 1750, and grew to over 1.6 million by 1850 (Tilastollinen päätoimisto,, 1899; Voutilainen et al.,, 2020). However, the healthcare system was primitive and virtually non-existent before the 19th century, lagging behind in development compared to the other Nordic countries and Europe until the 20th century (Saarivirta et al.,, 2009, 2012).
@@
Communication within Finland and between neighbouring countries, combined with limited knowledge of infections and a poor healthcare system, were factors that enabled recurring epidemics of various diseases, including measles. While armed conflicts between countries were not uncommon, wars and border changes affected, among many other things, the common welfare of people. Finland had been part of Sweden for centuries, but already in 1721, Russia annexed the southeastern parts of Finland, or Old Finland, including the significant trading center of Vyborg. After the Finnish War in 1808–-1809, the rest of Finland became an autonomous grand duchy of Russia. In 1812, Old Finland was united with the new autonomous Finland, and the capital of Finland was moved from Turku on the southwestern coast to Helsinki on the southern coast, impacting markets and trade connections (Ojala and Räihä,, 2017). Being a remarkable transformation, the administrative change, however, left also many things unchanged. For example, the demographic records collected during the Swedish administration were still maintained in a similar way as before.
@@
We analyse information obtained from parish records, using monthly data from January 1750 to December 1850 from $387$ different towns located in the southern half of Finland. The observations are dichotomous, indicating whether at least one death caused by measles was recorded. Thus, the data contains zeros for towns and months with no recorded deaths, and ones for those with recorded cases. Corresponding data, covering a different time span, has been used in Pasanen et al., (2023), which also describes the handling of the data in more detail. Other studies analysing data based on the same registers but covering different areas or time periods, or being aggregated to a different level, include, for example, Ketola et al., (2021) and Briga et al., (2021, 2022).
The data come with a significant amount of missing observations for several reasons. These include misdiagnoses, varying personal practices in maintaining the registers (Pitkänen,, 1977), and loss of records due to fires and other conditions over time. There are a total of $1212$ time points, each having observations from 75757575272272272272 sites simultaneously, so no month has information from all the sites at once. From another perspective, only one town has no missing observations at all, whereas all observations are missing from $44$ towns. Overall, at the beginning of the time series, the missingness is approximately 70%, decreasing steadily to about 20% at the end of the time series, so that 40% of all data are missing.
@@ 2.2 Hidden Markov Models
Discrete time hidden Markov models (HMM) (Rabiner,, 1989) are statistical models for temporal, or other sequentially ordered data, consisting of two essential components. There are, firstly, the hidden stochastic Markov process describing the unobserved, yet often interesting, part of the model, and secondly, the observed process dependent on the hidden component. With the word ”state”, often appearing in the terminology, we refer to some discrete, prevailing situations related to the phenomenon of interest, for example, two states could be the presence and absence of a disease. Next, we introduce the method in more detail.
@@
To perform fully Bayesian estimation of the HMM parameters and the corresponding latent states, we first run an MCMC targeting the marginal posterior of the model parameters, $p(\Omega,A,\rho|y)$, where $y=(y_{1},\ldots,y_{T})$. For this we need the marginal likelihood $p(y|\Omega,A,\rho)$, which can be computed with the forward part of the forward-backward algorithm (Baum and Petrie,, 1966; Rabiner,, 1989, and the references therein). The latent states $x=(x_{1},\ldots,x_{T})$ can then be sampled in the post-processing stage given the posterior samples of $\{\Omega,A,\rho\}$. This marginalisation approach commonly used with latent variable models allows much more efficient sampling than the approaches where the estimation is based on the joint posterior $p(x,\Omega,A,\rho)$, and is also strictly necessary when using efficient Hamiltonian Monte Carlo type of MCMC algorithms which rely on the differentiability of the posterior.
The marginal likelihood approach is also convenient with respect to missing data. When computing the emission probabilities $\omega_{s,t}$, i.e., the conditional probabilities of the observations given the current state and the model parameters, we set $\omega_{s,t}=1$ for all $s$ if the observation $y_{t}$ is missing. This means that as we do not gain any new information about the states due to the missing observation at time $t$, we essentially skip the contribution of $\omega_{s,t}$ at the marginal likelihood computation in forward algorithm at that time point (Zucchini and MacDonald,, 2009).
The likelihood and the posterior density of the hidden Markov models typically contain multiple modes. Some of these modes can be due to the label-switching problem (Jasra et al.,, 2005) but in complex models true local modes can also be present. In maximum likelihood setting the model parameters are often estimated by repeating the estimation several times with varying starting values for the optimisation algorithm, with the best solution declared as a global mode, while ignoring the other modes. In Bayesian setting multimodality of the posterior can pose both computational (convergence of the MCMC algorithms) and interpretability challenges. For example, it is difficult to interpret (label) and visualise the hidden states if the corresponding observational density is multimodal. These issues have to be taken into consideration when using HMMs.
The complexity of the response variables $y_{t}$ is not limited to scalars, but they can be multidimensional as well. In our case they are vectors containing spatially dependent elements. This means that the states, as well as the parameters, should allow for spatial structure.
@@ 2.3 Intrinsic conditional autoregressive models
Next we introduce the intrinsic conditional autoregressive (ICAR) model, which is a special case of the conditional autoregressive (CAR) model (Besag,, 1974), especially meant for spatial analysis to depict the amount of dependency between different locations. These models are based on a spatial division of a larger area, employing a neighbourhood definition for the generated subareas, and combining both of these with a Gaussian distribution in order to quantify the amount of the spatial dependency. This method does not offer information on the direction of the effects, i.e., the causal relationships, but only on the total amount of dependency.
@@
In case of ICAR model, we assume that $\alpha=1$, which results in improper distribution, which cannot be used as a model for the data, but as a prior distribution for spatial dependency, as is done for example in the Besag York Mollié (BYM) models (Besag et al.,, 1991). The parameters $\varphi_{i}$ are not identifiable without further constraints, as adding a constant to all of them does not change their density. A common way to avoid the identifiability problem is to restrict them to sum to zero over the sites.
Some reasons to employ the improper distribution of ICAR instead of the proper of the CAR, are the facts that, firstly, to account significant amount of spatial association, the parameter $\alpha$ most likely has to be close to one, and, secondly, the width of the posterior spatial pattern might be limited in the case of CAR (Banerjee et al.,, 2015). Thus, in some cases, it might be convenient to choose ICAR in the first place.
2.4 HMM with ICAR
In our approach, we combine the HMM and the spatial ICAR model, introduced above, thus enabling the HMM analysis of spatio-temporal data sets. We apply our method to the historical Finnish parish register data to study the dynamics of the measles epidemics. More precisely, our response variable $y_{i,t}$ is dichotomous, indicating the presence of at least one measles induced death observation in town $i$ in time point $t$.
As before, we denote each site with $i=1,\dots,N$, and each time point with $t=1,\dots,T$. In this setting, the probability of observing at least one death $y_{i,t}$ caused by measles in site $i$ and time point $t$ can be modeled as
@@ 2.4 HMM with ICAR
where the states
$x_{t}|x_{t-1}\sim\text{Categorical}(A_{x_{t-1},.}).$ (6)
The notation $A_{x_{t-1},.}$ means the row $x_{t-1}$ including all columns of the matrix $A$. As we are using our method to study the occurrences and absences of deaths, Bernoulli distribution is a natural choice for the model. With other kinds of response variables also the distribution should be changed, for example Poisson distribution for counts. In formula 5, which corresponds to the $p(y_{t}|x_{t})$ in Equation 2, the logit link function maps the combination of the explanatory terms onto a probability scale required by the Bernoulli distribution. Also the link function should be changed to an appropriate one when using a different distribution. The four explanatory terms describe the state specific constants, the local constants, the intensity of spatial interaction between the site and its neighbours, and a monthly seasonal variation. The formula 6 defines the connection between the underlying states $x_{t}$ and the transition matrix $A$.
The state specific constant $\mu_{x_{t}=s}$, $s\in\{1,\dots,S\}$, sets a base level for the nationwide probability of observing at least one death in each state $s$. In order to reduce the common, yet problematic, multimodality encountered with HMMs, we define the state specific constants so that they are monotonically increasing in a similar way as described in Bürkner and Charpentier, (2020). Due to this, our states are ordered based on the general incidence level. In our case, to begin with, we define the constants for the first and last state, $\mu_{1}$ and $\mu_{S}$, respectively. We use these two to scale the ones for the other states. We do this by defining a simplex $\boldsymbol{m}$ of S components, i.e., a vector summing to one, with the first term $m_{1}$ fixed to zero, and defining the rest of the constants as
@@
To allow local deviations from the nationwide base level, we add a site dependent constant $\lambda_{i}$, which in turn is constant over the states. As we set the mean of the local constants to zero to improve model identifiability and MCMC sampling efficiency, we also gain a more straightforward interpretation of the state dependent constant $\mu_{s}$ as a nationwide average level. The local terms $\lambda_{i}$ are not spatially dependent by definition, but aim to capture any heterogeneous features of the sites, for example, differing population sizes or communication intensities in our case.
The actual spatial dependence between the sites is introduced in the third term, $\sigma_{\varphi}\varphi_{i,x_{t}}$. The dependency is achieved via the ICAR structure outlined above, i.e., adapting the formulas 3 and 4, and letting $\boldsymbol{\varphi}_{x_{t}=s}\sim\text{N}(0,(D-W)^{-1})$, where $s\in\{1,\dots,S\}$, and $\sigma_{\varphi}=\tau^{-\frac{1}{2}}$. To employ the ICAR component, we need a definition for the neighbourhood. Here two regions are considered as neighbours if they share a border. Other definitions, for example, based on the distance from the centers of the sites, could be used as well. In practice, the matrices $W$ and $D$ are constructed according to our neighbourhood definition in order to access the covariance of the spatial terms. The state specific spatial term $\varphi_{i,x_{t}}$ is zero on average over the sites to guarantee identifiability using soft constraint as in Morris et al., (2019).
@@
As mentioned in subsection 2.1, our data are missing a large number of observations. However, as long as we assume they are missing at random, the HMM is a convenient approach to work with such imperfect data sets. Due to the hidden level, the observations do not depend on time, conditioned on the states, which is not the case when modelling dependencies only on the observed level. Now, the likelihood can be calculated without the missing observations by computing $p(y_{t}|x_{t})$ over only the observed sites at time $t$. Thus, the absent information can be omitted while estimating the model, yet the model provides estimates for all sites and time points.
@@
where $\text{N}(.,.)[z,]$ denotes a normal distribution truncated at $z$, and $\boldsymbol{m}_{2:S}$ denotes the components from $2$ to $S$ of the vector $\boldsymbol{m}$, omitting the first component since it is zero. The priors for the state specific constants $\mu_{1}$ and $\mu_{S}$ are chosen based on the data, to roughly match with the $10$th and $90$th percentiles of the monthly averages of the observed death occurrences. The row sums of the transition matrix $A$ are one, as well as the sums of the parameters $\rho$ and $a$, which is achieved with the Dirichlet priors. In our case, the number of time points is $T=1212$ and the number of sites $N=387$. For our final model, we set the number of states to be $S=5$.

### v2

@@ Abstract
Keywords: Bayes, epidemiology, hidden Markov model, intrinsic conditional autoregression, measles, spatio-temporal
@@ 1 Introduction
In this paper, we propose a Bayesian hidden Markov model with a state-dependent spatial correlation structure as an unsupervised learning method for decomposing the spatio-temporal variation in the data into relatively few interpretable latent components. We apply our model to study the spatio-temporal occurrence of measles in Finland during the 18th and 19th centuries, offering insights into the historical measles epidemics in Finland and the role of communication routes in disease transmission routes. Although effective measles vaccines have significantly reduced global incidence, measles remains a major public health concern, as evidenced by ongoing outbreaks in various parts of the world. In addition to its historical significance, Finland in the 18th and 19th centuries illustrates how measles spread in sparsely populated areas with limited healthcare access, conditions still present in some contemporary settings. Overall, this type of descriptive analysis in epidemiology may be of great value (Fox et al., 2022), and historical settings such as ours allow the study of epidemic patterns under more natural conditions uninfluenced by modern healthcare, thus contributing to the basic biological understanding of measles.
@@
Describing, visualising, and modelling spatio-temporal panel data is often challenging, particularly in high-dimensional settings with numerous time points and spatial areas. Conditional autoregressive models (CAR) (Besag, 1974) and their variants, such as intrinsic conditional autoregressive models (ICAR), are frequently employed to describe areal data. Although it is possible to extend CAR models to spatio-temporal settings (Clayton and Bernardinelli, 1996), these interaction models can become computationally intensive with high-dimensional data where spatial and temporal effects are not separable (Knorr-Held, 2000). Approximate methods like INLA (Rue et al., 2009) may be computationally efficient alternatives to asymptotically exact Bayesian estimation via Markov chain Monte Carlo (MCMC), but these methods typically impose some restrictions on the model structure and the number of hyperparameters (Bakka et al., 2018). The potential bias of these approximations is also challenging to quantify in specific cases, although in many typical settings the approximation bias has been found to be negligible (Ogden, 2017; Rue et al., 2017). Furthermore, visualising and interpreting the results of these interaction models may be complex: with time-varying spatial effects and a large number of time points, a substantial number of static maps are required to convey information about the estimated (or observed) spatial dynamics.
@@
In life course research, hidden Markov models have been suggested as a probabilistic method to condense information of complex, multidomain life courses into a few interpretable latent states (Helske et al., 2018; Helske and Helske, 2019). In these applications, given a common data generating process, individuals are assumed to be independent, having their own latent trajectories. In contrast, in spatio-temporal panel data, individuals share a spatial dependency structure which should be accounted for.
@@
In Amorós et al. (2020), hidden Markov models are utilised in a spatio-temporal setting to detect influenza outbreaks, and in Knorr-Held and Richardson (2003), they are used to identify meningococcal disease incidences. Both studies use two states to identify the endemic and epidemic, or endemic and hyperendemic, phases of the infections but do not aim to deeply analyse the underlying dynamics of the diseases. Employing more states may aid in discovering latent phenomena related to the epidemics. Amorós et al. (2020) include the spatial dependence in the epidemic state as temporally independent ICAR components and omit it in the endemic state, whereas Knorr-Held and Richardson (2003) account for the spatial dependency as ICAR in both states but consider it as state and time-invariant.
@@
Extremely low population size coupled with hundreds of sparsely spaced small towns makes it challenging for epidemics to remain fully endemic (Keeling and Grenfell, 1997). Thus, it is probable that in such conditions of Finland, fadeouts and reintroductions dominated the infection dynamics, implying that changes in the transmission sources could have strongly affected these dynamics. Finland and Sweden are separated by a large geographical barrier, the Gulf of Bothnia, and this, as well as the Russian border on the east, must have influenced the trade connections and the movement of people. The annexation of Finland to Russia, after a war in 1808–1809, resulted in the transfer of the capital from the western position of Turku to a more eastern location of Helsinki, increased dramatically trade and social communication, and altered drastically Finland’s connectivity to other countries (Ojala and Räihä, 2017). We exploit our method to describe the reoccurring and presumably significantly foreign connection driven patterns of the epidemics in a compact yet informative way. [^1] This text was originally published as a preprint (https://arxiv.org/abs/2405.16885v1).
@@ 2 Data and methods
In this section, first, we describe the data and their background in a historical and Finnish context. The rest of the chapter progresses from presenting the model concepts of the statistical components needed for our approach to presenting the main model itself. The readers familiar with the hidden Markov models and intrinsic conditional autoregression or with less interest in the theoretical details of those may also give less attention to the next subsections 2.2 and 2.3 and continue reading straight from the subsection 2.4. In that we represent in detail the main model and the necessary technical assumptions to estimate it, recalling the connection between the practice, given our application, and the statistical theory.
@@ 2.1 Data
Finland has a long history in collecting demographic and health-related data, with local parishes systematically maintaining records of baptisms, burials, and causes of death from 1749 onwards. Also earlier, albeit incomplete, records exist (Pitkänen, 1977). Even though death diagnostics were symptom-based, the distinctive features of certain diseases allowed for relatively accurate diagnoses. One of those infections is measles, which we focus on. According to our data, measles accounted for approximately 2% of all deaths during our study period from 1750 to 1850. The population of Finland was roughly $500{,}000$ in 1750, and grew to over 1.6 million by 1850 (Tilastollinen päätoimisto, 1899; Voutilainen et al., 2020). However, the healthcare system was primitive and virtually non-existent before the 19th century, lagging behind in development compared to the other Nordic countries and Europe until the 20th century (Saarivirta et al., 2009, 2012).
@@
Communication within Finland and between neighbouring countries, combined with limited knowledge of infections and a poor healthcare system, were factors that enabled recurring epidemics of various diseases, including measles. While armed conflicts between countries were not uncommon, wars and border changes affected, among many other things, the common welfare of people. Finland had been part of Sweden for centuries, but in 1721, Russia annexed the southeastern parts of Finland, or Old Finland, including the significant trading center of Vyborg. After the Finnish War in 1808–-1809, the rest of Finland became an autonomous grand duchy of Russia. In 1812, Old Finland was united with the new autonomous Finland, and the capital of Finland was moved from Turku on the southwestern coast to Helsinki on the southern coast, impacting markets and trade connections (Ojala and Räihä, 2017). Being a remarkable transformation, the administrative change, however, left also many things unchanged. For example, the demographic records collected during the Swedish administration were still maintained in a similar way as before.
We analyze deaths from measles recorded in the parish registers, using monthly data from January 1750 to December 1850 from $387$ different towns located in the southern half of Finland. The original records have count information on the deaths at the town level, but since the counts are generally low (99% of all measles counts in our data are three or fewer) and there are no town-level population sizes available—only nationwide estimates (Voutilainen et al., 2020)—it is challenging to separate the actual differences between towns from those resulting from the varying population sizes.
@@
To overcome these issues, we summarise the data into a more robust form. The observations we use are dichotomous, indicating whether at least one death caused by measles was recorded in each town. Thus, the data contains zeros for towns and months with no recorded deaths, and ones for those with recorded cases. Corresponding data, covering a different time span, has been used in Pasanen et al. (2024), which also describes the handling of the data in more detail. Other studies analysing data based on the same registers but covering different areas or time periods, or being aggregated to a different level, include, for example, Ketola et al. (2021) and Briga et al. (2021, 2022). Pasanen et al. (2024) also implies that the dichotomisation of the observations results in conclusions corresponding to those of undichotomised data.
Although in general the cause of death records contain errors due to misdiagnoses, measles was likely diagnosed accurately because of its distinct visible symptoms. However, the dataset exhibits substantial missingness of records attributable to heterogeneity in record-keeping practices across towns (Pitkänen, 1977) and to the destruction of records from fires and other historical events. Therefore, the missingness does not depend on the outcome variable itself. The study period spans $1212$ months, each having observations from 75757575272272272272 towns per month, so no month has complete information from all towns. Only one town has complete records, whereas all observations are missing from $44$ towns. The missingness is approximately 70% at the beginning of the study period, decreasing steadily to about 20% at the end of the time series, with an overall missingness rate approximately 40%. Based on these patterns, we assume that the response variable is missing at random (MAR), conditional on time and town.
In order to present our solution to analysing these data, next, we introduce the concepts of hidden Markov models and intrinsic conditional autoregression. The former compresses the time dependency into a relatively low number of nationwide discrete infection statuses, whereas the latter is responsible for considering the spatial connections.
@@ 2.2 Hidden Markov Models
Discrete time hidden Markov models (HMM) (Rabiner, 1989) are statistical models for temporal, or other sequentially ordered data, consisting of two essential components. There are, firstly, the hidden stochastic Markov process describing the unobserved, yet often interesting, part of the model, and secondly, the observed process dependent on the hidden component. With the word ”state”, often appearing in the terminology, we refer to some discrete, prevailing situations related to the phenomenon of interest. For example, in our case, there are two states indicating the presence and absence of a disease, and the observed process consists of the actual observations of the deaths caused by that disease. Next, we introduce the method in more detail.
@@
To perform a fully Bayesian estimation of the HMM parameters and the corresponding latent states, we first run an MCMC targeting the marginal posterior of the model parameters, $p(\Omega,A,\rho|y)$, where $y=(y_{1},\ldots,y_{T})$. For this we need the marginal likelihood $p(y|\Omega,A,\rho)$, which can be computed with the forward part of the forward-backward algorithm (Baum and Petrie, 1966; Rabiner, 1989, and the references therein). The latent states $x=(x_{1},\ldots,x_{T})$ can then be sampled in the post-processing stage given the posterior samples of $\{\Omega,A,\rho\}$. This marginalisation approach commonly used with latent variable models allows much more efficient sampling than the approaches where the estimation is based on the joint posterior $p(x,\Omega,A,\rho)$, and is also strictly necessary when using efficient Hamiltonian Monte Carlo type of MCMC algorithms which rely on the differentiability of the posterior.
The marginal likelihood approach is also convenient with respect to missing data under MAR. When computing the emission probabilities $\omega_{s,t}$, i.e., the conditional probabilities of the observations given the current state and the model parameters, we set $\omega_{s,t}=1$ for all $s$ if the observation $y_{t}$ is missing. This means that as we do not gain any new information about the states due to the missing observation at time $t$, we essentially skip the contribution of $\omega_{s,t}$ at the marginal likelihood computation in forward algorithm at that time point (Zucchini and MacDonald, 2009).
The likelihood and the posterior density of the hidden Markov models typically contain multiple modes. Some of these modes can be due to the label-switching problem (Jasra et al., 2005) but in complex models true local modes can also be present. In maximum likelihood setting the model parameters are often estimated by repeating the estimation several times with varying starting values for the optimisation algorithm, with the best solution declared as a global mode, while ignoring the other modes. In Bayesian setting multimodality of the posterior can pose both computational (convergence of the MCMC algorithms) and interpretability challenges. For example, it is difficult to interpret (label) and visualise the hidden states if the corresponding observational density is multimodal. These issues have to be taken into consideration when using HMMs.
The complexity of the response variables $y_{t}$ is not limited to scalars, but they can be multidimensional as well. In our case, they are vectors containing spatially dependent elements. This means that the states, as well as the parameters, should allow for spatial structure.
@@ 2.3 Intrinsic conditional autoregressive models
Next, we introduce the intrinsic conditional autoregressive (ICAR) model, which is a special case of the conditional autoregressive (CAR) model (Besag, 1974), especially meant for spatial analysis to depict the amount of dependency between different locations. These models are based on a spatial division of a larger area, employing a neighbourhood definition for the generated subareas, and combining both of these with a Gaussian distribution in order to quantify the amount of the spatial dependency. This method does not offer information on the direction of the effects, i.e., the causal relationships, but only on the total amount of dependency.
@@
The normal distribution of formula 3 can also be written as
$p(\boldsymbol{\varphi})=(2\pi)^{N/2}\left|(\tau(D-\alpha W))^{-1}\right|^{1/2}\text{exp}\left(-\frac{\tau}{2}\boldsymbol{\varphi}^{\top}(D-\alpha W)\boldsymbol{\varphi}\right)\!,$ (5)
where $\boldsymbol{\varphi}^{\top}$ denotes the transpose of $\boldsymbol{\varphi}$ and $|x|$ stands for the determinant of $x$. In case of ICAR model, we assume that $\alpha=1$, which results in improper distribution, which cannot be used as a model for the data, but as a prior distribution for spatial dependency, as is done for example in the Besag York Mollié (BYM) models (Besag et al., 1991). The parameters $\varphi_{i}$ are not identifiable without further constraints, as adding a constant to all of them does not change their density. A common way to avoid the identifiability problem is to restrict them to sum to zero over the sites.
Some reasons to employ the improper distribution of ICAR instead of the proper of the CAR, are the facts that, firstly, to account significant amount of spatial association, the parameter $\alpha$ most likely has to be close to one, and, secondly, the width of the posterior spatial pattern might be limited in the case of CAR (Banerjee et al., 2015). Additionally, when employing the ICAR model, the specification $\alpha=1$ obviates the need to calculate the determinant in Equation 5 since it becomes constant, thus reducing the computation time even from days to hours (Morris et al., 2019). Therefore, in some cases, it might be convenient to choose ICAR in the first place.
2.4 Main model: HMM with ICAR
In our approach, we combine the HMM and the spatial ICAR model, introduced above, thus enabling the HMM analysis of spatio-temporal data sets. We apply our method to the previously introduced historical Finnish parish register data to study the dynamics of the Finnish measles epidemics in 1750–1850. More precisely, we model the probability of observing at least one measles induced death in each town at each month.
We denote each site, i.e. town, with $i=1,\dots,N$, and each time point, i.e. month, with $t=1,\dots,T$. In this setting, the probability of observing at least one death $y_{i,t}$ caused by measles in site $i$ and time point $t$ can be modeled as
@@ 2.4 Main model: HMM with ICAR
where the latent, discrete states
$x_{t}|x_{t-1}\sim\text{Categorical}(A_{x_{t-1},.})$ (7)
represent the prevailing nationwide infection statuses. The notation $A_{x_{t-1},.}$ means the row $x_{t-1}$ including all columns of the transition matrix $A$. As we are using our method to study the occurrences and absences of deaths, Bernoulli distribution is a natural choice for the model. With other kinds of response variables also the distribution should be changed, for example Poisson distribution for counts. In formula 6, which corresponds to the $p(y_{t}|x_{t})$ in Equation 2, the logit link function maps the combination of the explanatory terms onto a probability scale required by the Bernoulli distribution. Also the link function should be changed to an appropriate one when using a different distribution. The four explanatory terms describe the state specific constants, the local constants, the intensity of spatial interaction between the site and its neighbours, and a monthly seasonal variation. The formula 7 defines the connection between the underlying states $x_{t}$ and the transition matrix $A$.
The state specific constant $\mu_{x_{t}=s}$, $s\in\{1,\dots,S\}$, sets a base level for the nationwide probability of observing at least one death in each state $s$. In order to reduce the common, yet problematic, multimodality encountered with HMMs, we define the state specific constants so that they are monotonically increasing in a similar way as described in Bürkner and Charpentier (2020). Due to this, our states are ordered based on the general incidence level. In our case, to begin with, we define the constants for the first and last state, $\mu_{1}$ and $\mu_{S}$, respectively. We use these two to scale the ones for the other states. We do this by defining a simplex $\boldsymbol{m}$ of S components, i.e., a vector summing to one, with the first term $m_{1}$ fixed to zero, and defining the rest of the constants as
@@
To allow local deviations from the nationwide base level, we add a site dependent term $\lambda_{i}$, which is constant over the states, i.e., over time. These local terms do not have a spatial dependency structure by definition, but they aim to capture any heterogeneous, time constant features of the sites, for example, differing population densities or communication intensities in our case. As we set the mean of the local constants to zero to improve model identifiability and MCMC sampling efficiency, we also gain a more straightforward interpretation of the state dependent constant $\mu_{s}$ as a nationwide average level.
The actual spatial dependence between the sites is introduced in the third term, $\sigma_{\varphi}\varphi_{i,x_{t}}$. The dependency is achieved via the ICAR structure outlined above, i.e., adapting the formulas 3 and 4, and letting $\boldsymbol{\varphi}_{x_{t}=s}\sim\text{N}(0,(D-W)^{-1})$, where $s\in\{1,\dots,S\}$, and $\sigma_{\varphi}=\tau^{-\frac{1}{2}}$. To employ the ICAR component, we need a definition for the neighbourhood. First of all, our sites are the towns based on the parishes, i.e., they are areas with borders. Here, two regions are considered as neighbours if they share a border. Other definitions, for example, based on the distance from the centers of the sites, could be used as well. In practice, the matrices $W$ and $D$ are constructed according to our neighbourhood definition in order to access the covariance of the spatial terms. The state specific spatial term $\varphi_{i,x_{t}}$ is zero on average over the sites to guarantee identifiability using soft constraint as in Morris et al. (2019).
@@
As mentioned in subsection 2.1, our data are missing a large number of observations. However, as long as we assume they are missing at random, the HMM is a convenient approach to work with such imperfect data sets. Due to the hidden level, the observations do not depend on time, conditioned on the states, which is not the case when modelling dependencies only on the observed level. Now, the likelihood can be calculated without the missing observations by computing $p(y_{t}|x_{t})$ over only the observed sites at time $t$ (Zucchini and MacDonald, 2009). Thus, missing observations are naturally handled during model estimation, yet the model provides estimates for all sites and time points, with the missing cases influencing the posterior uncertainty of the estimates.
@@
where $\text{N}(.,.)[z,]$ denotes a normal distribution truncated at $z$, and $\boldsymbol{m}_{2:S}$ denotes the components from $2$ to $S$ of the vector $\boldsymbol{m}$, omitting the first component since it is zero. The priors for the state specific constants $\mu_{1}$ and $\mu_{S}$ are chosen based on the data, to roughly match with the $10$th and $90$th percentiles of the monthly averages of the observed death occurrences. Such utilisation of the data is a common practice, for example, in rstanarm (Muth et al., 2018). The row sums of the transition matrix $A$ are one, as well as the sums of the parameters $\rho$ and $a$, which is achieved with the Dirichlet priors. These priors are mostly weakly informative and chosen to enhance the computational efficiency (Muth et al., 2018; Banner et al., 2020). In our case, the number of time points is $T=1212$ and the number of sites $N=387$. For our final model, we set the number of states to be $S=5$.